Find the perimeter of a rectangle given below:
__________________
| |
| |
| A= 250 sq ft | |
| |
|__________________|
25 ft

Answer by Evy(14)

Hello, I think i can help you! lol
Well first we need to find the other side of the rectangle in order to get the perimeter. This is simply done by dividing.You Start with the formula A=L*W.


so your equation is, 250=L*25


Now divide 250 by 25.
you are now left with 10=L. So the length is 10.


The formula for the perimeter is P=2(L+W).
Now we can plug in the numbers.


P= 2(10+25)


Add 10 and 25 Which gives you 35
You are now left with: P=2(35)


Multiply 35 by 2 the answer is 70

SO Perimeter=70 ft.


If you have any questions you can e-mail me at Southqueensb@yahoo.com
~EVY~